Determine if the following ratios form a proportion. Also, write the middle terms and extreme terms where the ratios form a proportion. 
2 kg : 80 kg and 25 g : 625 g 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2 kg : 80 kg and 25 g : 625 g
2 / 80 = 1 / 40
25 / 625 = 1 / 25
No, these are not in a proportion

Thanks